OpenStudy (anonymous):

the set (-1,1) is A.continuous B.discrete C.finite D.infinite

OpenStudy (anonymous):

discrete

OpenStudy (anonymous):

can u help me with a few more

OpenStudy (anonymous):

sure if i can

OpenStudy (anonymous):

let f and g be two functions defined by f(x)=x+5 with the domain all real numbers g(x)=2x,with domain all real numbers which of the following statements are true for f and g check all that apply A.the domains f and g are equal B.the ranges f and g are equal C.the functions f and g are equal

OpenStudy (anonymous):

hold up im thinking

OpenStudy (anonymous):

@issaiahphillips

OpenStudy (anonymous):

@maguihajji u have it yet??

OpenStudy (anonymous):

no im sorry ;'( i didnt really understand it in school either

OpenStudy (anonymous):

can u help me with more like the very first question

OpenStudy (anonymous):

construct a set in the following way .start with the set (1,2) .add all other positive powers of 2 to the set (that is 4,8,16 and so on ) until you get a power of 2 larger than 50 this set is A.continuous B.infinite C.finite D.discrete

OpenStudy (anonymous):

would it be continuous ?

OpenStudy (anonymous):

no no it cant be , its finite and discrete ( they mean the same thing btw )

OpenStudy (anonymous):

the set (0,-2,2,-4,4,-6,6) is a.continuous b.infinite c.finite d.discrete

OpenStudy (anonymous):

not really sure about that HELP @issaiahphillips

OpenStudy (anonymous):

@issaiahphillips

OpenStudy (anonymous):

lol here i am

OpenStudy (anonymous):

can u help with the question

OpenStudy (anonymous):

yeah it is B

OpenStudy (anonymous):

thats the first one the second is

OpenStudy (anonymous):

c or d

OpenStudy (anonymous):

let f and g be two functions defined by f(x)=x-1with the domain all even integers g(x)=x+1with the domain all even integers A.the domains f and g are equa B.the functions f and g are equal C.the ranges f and g are equal

OpenStudy (anonymous):

its c because the diffrence every time the go up or down is thwe same

OpenStudy (anonymous):

thanks for the help